# QNPU: Quantum Network Processor Unit for Quantum Supercomputers

### v2

@@ 4. Programming Model for Distributed Quantum Computing
While distributed quantum computing offers significant advantages for scaling quantum applications, it introduces complexity in programming and EPR resource management. In this section, we introduce a programming model which extends the OpenQASM (Cross et al., 2022) to simplify the development of distributed quantum algorithms.
@@ 4.1. DistQASM: Extending OpenQASM for Distributed Quantum Computing
We propose DistQASM, an extension of OpenQASM (Cross et al., 2022) designed specifically for distributed quantum computing. DistQASM builds upon OpenQASM’s human-readable representation of quantum circuits while adding critical features for distributed quantum computing. By extending rather than replacing OpenQASM, DistQASM maintains compatibility with the existing quantum software ecosystem while enabling the expression of distributed quantum operations. We introduce several key extensions in DistQASM:
@@
Direct Protocol Specification to define instructions for quantum communication protocols so that user can directly choose the protocol when writing the distributed quantum program. For the TP-Comm protocol, we define the teleport instruction to teleport a state from local node to remote node. For the CAT-COMM protocol, we define two DistQASM instructions: cat_ent and cat_disent. The cat_ent instruction will entangle the qubit in local node with the qubit in remote node. The cat_disent will disentangle these two qubits.
@@ 4.2. DistQASM Instruction Translation
Communication qubit allocation and zone transition: An available qubit is requested from the communication zone, and the qubit state is transitioned from the computation zone to the communication zone. This step reflects our decoupled QPU architecture, where computation and communication resources are physically separated. We assume that the QPU ISA provides instructions to perform this zone transition.
@@ 5. QNPU ISA Design
Abstraction of Quantum Communication Protocols: Instructions related to communication protocols initially flow from the QPU to the QNPU to execution, and QPU does not need to know the implementation details of the communication protocol. Therefore, our ISA should provides concise, well-defined protocol instructions that enable efficient instruction transfer between these components while abstracting the underlying complexity of quantum communication.
@@ 5.1. Communication Protocol Instructions
GET_TP_QUBIT <qubit_reg>: This instruction is a complementary instruction with SEND_TP, which handles the operations in the destination node. It identifies the teleported qubit and saves its index into the specified register, enabling the remote QPU to access the teleported state.register.
SEND_CAT_ENT_QUBIT <qubit_reg>, <dst_node_id>: Executes the CAT-Comm protocol to entangle a quantum state from a source qubit in the source node to the destination node. This instruction handles the operations in the source node.
GET_CAT_ENT_QUBIT <qubit_reg>: This instruction is a complementary instruction with SEND_CAT_ENT, which handles the operations in the destination node. It identifies the entangled qubit and saves its index into the specified register for subsequent operations by the remote QPU.
SEND_CAT_DISENT_QUBIT <qubit_reg>, <src_node_id>: Executes the disentanglement part of the CAT-Comm protocol to dis-entangle a qubit from the destination node to the source node.
GET_CAT_DISENT_QUBIT <qubit_reg>: This instruction is a complementary instruction with SEND_CAT_DISENT.
@@ 5.2.1. EPR Resource Management ops
EPR_RESERVE <EPRIdReg>, <dst_node_id>: The local node’s QNPU executes this op to occupy a prefetched EPR pair for quantum communication by checking the EPR resource table. If an available EPR pair is found, its ID is saved into the register EPRIdReg.
EPR_RESERVE_SYNC <StatusReg>: The remote node’s QNPU executes this op to synchronize EPR information with the local node, with the synchronization status saved in StatusReg. This ensures resource consistency across the distributed quantum nodes.
@@ 5.2.3. Quantum ops
Quantum ops implement the local quantum operations required for communication protocols. These include standard quantum gates and measurements as described in previous works (Fu et al., 2017, 2019). With our execution mdoel, these ops are executed by the QNPU specifically for communication-related quantum operations, while the QPU handles computation-zone quantum operations.
5.2.4. Example: Micro-operation Sequences for Communication Protocols
To illustrate how the communication protocol instructions are decoded into ops, we present the op sequences for the TP-Comm protocol related instructions, which includes the two complementary instruction SEND_TP_QUBIT and GET_TP_QUBIT.
@@ 5.2.4. Example: Micro-operation Sequences for Communication Protocols
In Listing 3, EPR_RESERVE, GET_EPR_QUBIT, and EPR_RELEASE are EPR management ops; SEND_EPR_ID, ACK_WAIT, and
TP_SEND_BITS are classical communication ops; and CNOT, H, MEAS are quantum ops. They are colored coded in the listings.
@@ 6.1. QNPU Scalar Pipeline Design
The QNPU serves as the communication interface between quantum nodes, enabling the execution of quantum communication protocols. Similar to the QPU, the QNPU consists of three components: (1) a classical controller, (2) a qubit control and readout interface for the communication qubits in the communication zone and (3) communication qubits in the communication zone. The QNPU classical controller receives the quantum communication protocol instructions from its QPU and decodes them into sequences of ops. These ops are then executed by specialized execution units and ultimately translated into control pulses for the qubits in the communication zone.
While previous works (Fu et al., 2017, 2019; Zhang et al., 2021) have discussed QPU microarchitecture design, this paper presents a novel QNPU microarchitecture specifically designed for distributed quantum computing. Our proposed pipeline architecture, shown in Figure 6, efficiently processes the communication protocol instructions. The key components of our QNPU microarchitecture include:
@@
Quantum ops Execution Unit: Executes quantum operations on the communication qubits.
@@
In order to show how the QNPU instructions run on the QNPU, we now present how the QNPU for a source Node A and a destination Node B execute the complementary instruction SEND_TP_QUBIT and GET_TP_QUBIT respectively. The instructions for the Cat-Comm protocol execute in a similar way.
@@ 6.2. QNPU Superscalar Design
As shown in Figure 7, our superscalar QNPU architecture incorporates multiple Decoder where each decoder independently processes communication protocol instructions, allowing simultaneous decoding of multiple instructions. Each decoder has its own ops buffer and execution units, preventing resource contention between parallel instruction streams. This architecture offers significant advantages for distributed quantum computing. When one decoder stalls due to outstanding ops in its buffer, incoming communication protocol instructions can be routed to available decoders. The QPU’s dependency checking ensures that only instructions that can safely execute in parallel are forwarded to available decoders, maintaining computational correctness while exploiting available parallelism.
This superscalar design is particularly valuable for executing distributed quantum circuits containing multiple independent remote quantum operations. By processing multiple communication protocol instructions in parallel, the superscalar QNPU significantly reduces the overall execution time, accelerating distributed quantum algorithms.
@@ 7.1. Experiment Setup
Performance Simulator: We developed a cycle-level simulator to evaluate the timing behavior of our proposed decoupled architecture. The simulator can be configured to model both the scalar and superscalar QNPU designs. We assume the availability of a compiler that efficiently schedules EPR pair generation, as demonstrated in prior work (Zhang et al., 2025). With such compiler support, we model perfect EPR scheduling and pre-generation, which hides the latency of EPR generation.
Baseline: We implement an in-order pipelined scalar QPU that executes instructions for both local and remote gates. In other words, it supports pipelined execution but not parallel execution of instructions. For fairness, we assume this monolithic QPU also has the same compiler support as our decoupled architecture, i.e., perfect EPR scheduling and pre-generation.
Benchmarks: Our benchmarks include Hamiltonian Simulation, Greenberger–Horne–Zeilinger (GHZ), Bernstein-Vazirani (BV), Quantum Fourier Transform (QFT), Variational Quantum Eigensolver (VQE), and Quantum Approximate Optimization Algorithm (QAOA). For Hamiltonian simulation, we study one-dimensional Transverse Field Ising Models (TFIM) of varying sizes. For VQE, we adopt the two-local ansatz from the Qiskit circuit library, with full entanglement layers or linear entanglement layers. For QAOA, we evaluate MaxCut problems on randomly generated graphs, where half of all possible edges are connected.
@@ 7.2. Experiment Results
We evaluate the execution time of distributed quantum workloads under three architectural configurations: (1) a monolithic QPU architecture (i.e., baseline), (2) a decoupled design with a scalar QNPU, and (3) a decoupled design with a 4-way superscalar QNPU. Table 2 summarizes the results of benchmarks with different circuit sizes and node counts.
Decoupled vs. Monolithic Architectures: Across all benchmarks, the decoupled architecture with a scalar QNPU provides only marginal improvements over the monolithic baseline, typically below 1%. This indicates that simply removing contention between computation and communication yields limited benefits, as the execution time of local gates is small compared to that of remote gates. In contrast, the decoupled architecture with a superscalar QNPU achieves substantial speedups, often exceeding 70%, for communication intensive benchmarks like QFT, VQE, and QAOA. These results show that the true advantage of our proposed decoupling architecture lies in separating computation from communication, which simplifies the design of superscalar communication pipelines and enables efficient parallel execution of remote gates.
Circuit-size Scaling: In this experiment, we increase the number of qubits on each node by increasing the circuit size while keeping the number of nodes the same. For Hamiltonian simulation benchmarks, the max number of remote gates per node remains constant as the circuit size grows (e.g., from 50 to 150 qubits, the max number of remote gates per node remains 4), and these remote operations are strictly serialized. Thus, both scalar and superscalar QNPUs yield negligible gains compared to the monolithic baseline. In contrast, for QFT, VQE, and QAOA circuits, when the circuit size increases from 50 to 150 qubits, both the total and per-node number of remote gates increase. In these cases, the superscalar QNPU delivers high reductions in execution time (70–75%), as wider pipelines effectively exploit parallelism across independent remote gates.
Node-count Scaling: In this experiment, we evaluate scaling of the number of nodes while fixing the circuit size to 150 qubits. For Hamiltonian simulation, increasing the node count does not increase the maximum number of remote gates per node. However, because all remote operations across nodes are serialized, distributing the circuit over more nodes introduces additional inter-node dependencies, which increases the overall execution time. In contrast, for benchmarks such as QFT, VQE, and QAOA, distributing the same circuit across more nodes increases the total number of remote gates but decreases the maximum number of remote gates assigned to each node. This reduction in per-node communication burden shortens execution time, leading all three architectures to benefit from lower latency as node count increases.
@@ 7.3. Impact of Superscalar Width
In the previous subsection 7.2, we used a fixed 4-way superscalar QNPU. To better understand how the pipeline width influences performance, we extend our study to multiple configurations (2-, 4-, 6-, 8-, 10-, 12-, 14-, and 16-way) and analyze their impact across different benchmarks. We focus on QFT, VQE, and QAOA, since these workloads contain remote operations that can potentially be executed in parallel. QFT circuits involve all-to-all interactions via controlled-phase gates, making them highly communication-intensive when distributed across nodes. VQE circuits with a full-entanglement ansatz also exhibit communication-intensive behavior. In contrast, our QAOA evaluation targets MaxCut on 2-regular graphs, producing relatively sparse circuits with fewer parallelizable remote operations than QFT and VQE.
Circuit-size scaling: Figure 8 shows results for QFT, VQE, and QAOA circuits of 30, 60, and 90 qubits distributed across 5 nodes. With a fixed node count, increasing circuit size introduces more remote operations per node, creating greater opportunities for parallel execution. As a result, wider superscalar pipelines remain effective for larger circuits. For example, the 90-qubit QFT continues to benefit from up to 12- and even 16-way designs, whereas the 30-qubit QFT saturates at 6-way. Across benchmarks, the 90-qubit QAOA saturates around 10-way, while the 90-qubit QFT and VQE still show gains beyond 10-way. This difference arises from QAOA’s relatively sparse structure, which provides fewer independent remote gates to exploit in parallel.
Node-count scaling: Figure 9 reports results for 30-qubit QFT, VQE, and QAOA circuits distributed across 2, 5, 10, 15, and 30 nodes. For the fixed QFT-30-qubit circuit, the benefit of widening decreases as the node count increases. With 2 nodes, wider designs provide clear speedups; with 5 nodes, performance saturates at 6-way; with 10 nodes, at 4-way; and with 15 nodes, at 2-way. When distributed across 30 nodes, the superscalar design offers no benefit. This trend reflects the fact that distributing a fixed circuit over more nodes reduces the number of remote operations per node, thereby also reducing opportunities of independent remote gates that can be executed in parallel. VQE and QAOA exhibit similar patterns, though they saturate at different numbers of nodes.
@@ 8. Conclusions
In this paper, we present a novel quantum supercomputer architecture featuring decoupled processing units, with QPUs for computation and QNPUs for communication. We develop the instruction set architecture and microarchitecture for the QNPU. The implementation of our comprehensive instruction set architecture, along with scalar and superscalar QNPU designs, can effectively address key bottlenecks in distributed quantum computing, particularly for communication-intensive workloads. Furthermore, by extending OpenQASM to our proposed DistQASM for simplified programming of distributed quantum algorithms, we have provide a complete solution that bridges the gap between quantum applications and the underlying hardware infrastructure. To evaluate our design, we implement a cycle-level performance simulator that evaluates the timing behavior of the proposed architecture under distributed quantum workloads. The simulation results demonstrate that these architectural innovations deliver substantial performance improvements compared to conventional monolithic designs.

### v3

@@ 4. Programming Model for Distributed Quantum Computing
While distributed quantum computing offers significant advantages for scaling quantum applications, it introduces complexity in programming and EPR resource management. In this section, we introduce a programming model which extends OpenQASM (Cross2022QASM) to simplify the development of distributed quantum algorithms.
@@ 4.1. DistQASM: Extending OpenQASM for Distributed Quantum Computing
We propose DistQASM, an extension of OpenQASM (Cross2022QASM) designed specifically for distributed quantum computing. DistQASM builds upon OpenQASM’s human-readable representation of quantum circuits while adding critical features for distributed quantum computing. By extending rather than replacing OpenQASM, DistQASM maintains compatibility with the existing quantum software ecosystem while enabling the expression of distributed quantum operations. We introduce several key extensions in DistQASM:
@@
Direct Protocol Specification to define instructions for quantum communication protocols so that user can directly choose the protocol when writing the distributed quantum program. For the TP-Comm protocol, we define the teleport instruction to teleport a state from the source node to the destination node. For the CAT-COMM protocol, we define two DistQASM instructions: cat_ent and cat_disent. The cat_ent instruction entangles the qubit in the source node with the qubit in the destination node. The cat_disent disentangles these two qubits.
@@ 4.2. DistQASM Instruction Translation
Qubit allocation and zone transition: An available qubit is requested from the communication zone, and the qubit state is transitioned from the computation zone to the communication zone. This step reflects our decoupled QPU architecture, where computation and communication resources are physically separated. We assume that the QPU ISA provides instructions (e.g., SWAP or physical move) to perform this zone transition.
@@ 5. QNPU ISA Design
Abstraction of Quantum Communication Protocols: Instructions related to communication protocols initially flow from the QPU to the QNPU for execution, and QPU does not need to know the implementation details of the communication protocol. Therefore, our ISA should provides concise, well-defined protocol instructions that enable efficient instruction transfer between these components while abstracting the underlying complexity of quantum communication.
@@ 5.1. Communication Protocol Instructions
GET_TP_QUBIT <qubit_reg>: This instruction is a complementary instruction to SEND_TP, which handles the operations in the destination node. It identifies the teleported qubit and saves its index into the specified register, enabling the destination QPU to access the teleported state register.
SEND_CAT_ENT_QUBIT <qubit_reg>, <dst_node_id>: Executes the CAT-Comm protocol to entangle a quantum state from a qubit in the source node to the destination node. This instruction handles the operations in the source node.
GET_CAT_ENT_QUBIT <qubit_reg>: This instruction is a complementary instruction to SEND_CAT_ENT, which handles the operations in the destination node. It identifies the entangled qubit and saves its index into the specified register for subsequent operations by the destination QPU.
SEND_CAT_DISENT_QUBIT <qubit_reg>, <src_node_id>: Executes the disentanglement part of the CAT-Comm protocol to dis-entangle the qubit from the destination node to the source node.
GET_CAT_DISENT_QUBIT <qubit_reg>: This instruction is a complementary instruction to SEND_CAT_DISENT.
@@ 5.2.1. EPR Resource Management ops
EPR_RESERVE <EPRIdReg>, <dst_node_id>: The source node’s QNPU executes this op to occupy a prefetched EPR pair for quantum communication by checking the EPR resource table. If an available EPR pair is found, its ID is saved into the register EPRIdReg.
EPR_RESERVE_SYNC <StatusReg>: The destination node’s QNPU executes this op to synchronize EPR information with the source node, with the synchronization status saved in StatusReg. This ensures resource consistency across the distributed quantum nodes.
@@ 5.2.3. Quantum ops
Quantum ops implement the local quantum operations required for communication protocols. These include standard quantum gates and measurements as described in previous works (Fu2017QuMA; Fu2019eQASM). With our execution mdoel, these ops are executed by the QNPU specifically for communication-related quantum operations, while the QPU handles computation-zone quantum operations.
Example: To illustrate how the communication protocol instructions are decoded into ops, we present the op sequences for the TP-Comm protocol related instructions, which include the two complementary instructions, SEND_TP_QUBIT and GET_TP_QUBIT.
@@
In Listing 3, EPR_RESERVE, GET_EPR_QUBIT, and EPR_RELEASE are EPR management ops; SEND_EPR_ID, ACK_WAIT, and TP_SEND_BITS are classical communication ops; and CNOT, H, MEAS are quantum ops. They are colored coded in the listings.
@@ 6.1. QNPU Scalar Pipeline Design
The QNPU serves as the communication interface between quantum nodes, enabling the execution of quantum communication protocols. Similar to the QPU, the QNPU consists of three components: (1) a classical controller, (2) a qubit control and readout interface for the qubits in the communication zone and (3) qubits in the communication zone. The QNPU classical controller receives the quantum communication protocol instructions from its QPU and decodes them into sequences of ops. These ops are then executed by specialized execution units and ultimately translated into control pulses for the qubits in the communication zone.
While previous works (Fu2017QuMA; Fu2019eQASM; Zhang2021parallelArch) have discussed QPU microarchitecture design, this paper presents a novel QNPU microarchitecture specifically designed for distributed quantum computing. Our proposed pipeline architecture, shown in Fig. 6, efficiently processes the communication protocol instructions. The key components of our QNPU microarchitecture include:
@@
Quantum ops Execution Unit: Executes quantum operations on qubits in the communication zone.
@@
In order to show how the QNPU instructions run on the QNPU, we now present how the QNPUs for a source Node A and a destination Node B execute the complementary instructions, SEND_TP_QUBIT and GET_TP_QUBIT, respectively. The instructions for the Cat-Comm protocol execute in a similar way.
@@ 6.2. QNPU Superscalar Design
As shown in Fig. 7, our superscalar QNPU architecture incorporates multiple Decoder where each decoder independently processes communication protocol instructions, allowing simultaneous decoding of multiple instructions. Each decoder has its own ops buffer and execution units, preventing resource contention between parallel instruction streams. This architecture offers significant advantages for distributed quantum computing. When one decoder stalls due to outstanding ops in its buffer, incoming communication protocol instructions can be routed to available decoders.
This superscalar design is particularly useful for executing distributed quantum circuits containing multiple independent remote quantum operations. By processing multiple communication protocol instructions in parallel, the superscalar QNPU significantly reduces the overall execution time, accelerating distributed quantum algorithms.
@@ 7.1. Experiment Setup
Performance Simulator: We developed a cycle-level simulator to evaluate the timing behavior of our proposed decoupled architecture. The simulator can be configured to model both the scalar and superscalar QNPU designs. We assume the availability of a compiler that efficiently schedules EPR pair generation, as demonstrated in prior work (zhang2024QDC). With such compiler support, we model perfect EPR scheduling and pre-generation, which hides the latency of EPR generation.
Baseline: For comparison, we implement a baseline architecture following the unified processing model of NetQASM (Dahlberg2022NetQASM), where a single QNPU handles both local and remote quantum operations. Our baseline uses an in-order pipelined scalar QNPU that executes instructions for both local and remote gates. In other words, it supports pipelined execution but not parallel execution of instructions. For fairness, we assume this monolithic QNPU also has the same compiler support as our decoupled architecture, i.e., perfect EPR scheduling and pre-generation.
Benchmarks: Our benchmarks include Hamiltonian Simulation, Greenberger–Horne–Zeilinger (GHZ), Bernstein-Vazirani (BV), Quantum Fourier Transform (QFT), Variational Quantum Eigensolver (VQE), and Quantum Approximate Optimization Algorithm (QAOA). For Hamiltonian simulation, we study one-dimensional Transverse Field Ising Models (TFIM) of varying sizes. For VQE, we adopt the two-local ansatz from the Qiskit circuit library, with full entanglement layers or linear entanglement layers. For QAOA, we evaluate MaxCut problems on 2-regular graphs and randomly generated graphs. For randomly generated graphs, half of all possible edges are connected.
@@ 7.2. Experiment Results
We evaluate the execution time of distributed quantum workloads under three architectural configurations: (1) a monolithic QNPU architecture (i.e., baseline), (2) a decoupled design with a scalar QNPU, and (3) a decoupled design with a 4-way superscalar QNPU. Table 2 summarizes the results of benchmarks with different circuit sizes and node counts.
Decoupled vs. Monolithic Architectures: Across all benchmarks, the decoupled architecture with a scalar QNPU provides only marginal improvements over the monolithic baseline, typically around 1%. This indicates that simply removing contention between computation and communication yields limited benefits, as the execution time of local gates is small compared to that of remote gates. In contrast, the decoupled architecture with a superscalar QNPU achieves substantial speedups, often exceeding 70%, for communication intensive benchmarks like BV, QFT, VQE with full entanglement ansatz, and QAOA. These results show that the true advantage of our proposed decoupling architecture lies in separating computation from communication, which simplifies the design of superscalar communication pipelines and enables efficient parallel execution of remote gates.
Circuit-size Scaling: In this experiment, we increase the number of qubits on each node by increasing the circuit size while keeping the number of nodes the same. For Hamiltonian simulation, GHZ, and VQE with linear entanglement ansatz, the max number of remote gates per node remains constant as the circuit size grows (e.g., from 50 to 150 qubits, the max number of remote gates per node remains 4 for Hamiltonian simulation benchmarks), and these remote operations are strictly serialized. Thus, both scalar and superscalar QNPUs yield negligible gains compared to the monolithic baseline. In contrast, for BV, QFT, VQE with full entanglement ansatz, and QAOA circuits, when the circuit size increases from 50 to 150 qubits, both the total and per-node number of remote gates increase. In these cases, the superscalar QNPU delivers high reductions in execution time (70–75%), as wider pipelines effectively exploit parallelism across independent remote gates.
Node-count Scaling: In this experiment, we evaluate scaling of the number of nodes while fixing the circuit size to 150 qubits. For Hamiltonian simulation, GHZ, and VQE with linear entanglement ansatz, increasing the node count does not constantly increase the maximum number of remote gates per node. However, because all remote operations across nodes are serialized, distributing the circuit over more nodes introduces additional inter-node dependencies, which increases the overall execution time. In contrast, for benchmarks such as QFT, VQE with full entanglement ansatz, and QAOA, distributing the same circuit across more nodes increases the total number of remote gates but decreases the maximum number of remote gates assigned to each node. This reduction in per-node communication burden shortens execution time, leading all three architectures to benefit from lower latency as node count increases.
@@ 7.3. Impact of Superscalar Width
In the previous subsection 7.2, we used a fixed 4-way superscalar QNPU. To better understand how the superscalar width influences performance, we extend our study to multiple configurations (2-, 4-, 6-, 8-, 10-, 12-, 14-, and 16-way) and analyze their impact across different benchmarks. We focus on QFT, VQE with full entanglement ansatz, and QAOA, since these workloads contain remote operations that can potentially be executed in parallel. QFT circuits involve all-to-all interactions via controlled-phase gates, making them highly communication-intensive when distributed across nodes. VQE circuits with a full-entanglement ansatz also exhibit communication-intensive behavior. In contrast, our QAOA evaluation targets MaxCut on 2-regular graphs, producing relatively sparse circuits with fewer parallelizable remote operations than QFT and VQE.
Circuit-size scaling: Fig. 8 shows results for QFT, VQE, and QAOA circuits of 30, 60, and 90 qubits distributed across 5 nodes. With a fixed node count, increasing circuit size introduces more remote operations per node, creating greater opportunities for parallel execution. As a result, wider superscalar pipelines remain effective for larger circuits. For example, the 90-qubit QFT continues to benefit from up to 12- and even 16-way designs, whereas the 30-qubit QFT saturates at 6-way. Across benchmarks, the 90-qubit QAOA saturates around 10-way, while the 90-qubit QFT and VQE still show gains beyond 10-way. This difference arises from QAOA’s relatively sparse structure, which provides fewer independent remote gates to exploit in parallel.
Node-count scaling: Fig. 9 reports results for 30-qubit QFT, VQE, and QAOA circuits distributed across 2, 5, 10, 15, and 30 nodes. For the fixed QFT-30-qubit circuit, the benefit of larger superscalar width decreases as the node count increases. With 2 nodes, wider designs provide clear speedups; with 5 nodes, performance saturates at 6-way; with 10 nodes, at 4-way; and with 15 nodes, at 2-way. When distributed across 30 nodes, the superscalar design offers no benefit. This trend reflects the fact that distributing a fixed circuit over more nodes reduces the number of remote operations per node, thereby also reducing opportunities of independent remote gates that can be executed in parallel. VQE and QAOA exhibit similar patterns, though they saturate at different superscalar widths.
@@ 8. Conclusions
In this paper, we present a novel quantum supercomputer architecture featuring decoupled processing units, with QPUs for computation and QNPUs for communication. We develop the instruction set architecture and microarchitecture for the QNPU to improve communication efficiency particularly for communication-intensive workloads. Furthermore, by extending OpenQASM to our proposed DistQASM for simplified programming of distributed quantum algorithms, we have provide a complete solution that bridges the gap between quantum applications and the underlying hardware infrastructure. To evaluate our design, we implement a cycle-level performance simulator that evaluates the timing behavior of the proposed architecture under distributed quantum workloads. The simulation results demonstrate that these architectural innovations deliver substantial performance improvements compared to conventional monolithic designs.